OpenStudy (j.f):

What will the next number be? 1/2, 1/12, 1/30, ?

hero (hero):

Any ideas for this @J.F ?

hero (hero):

First know that we can break this down to: 1/(2*1), 1/(2*6), 1/(2*15) That's your first hint. Try to find a pattern with 1, 6, and 15

OpenStudy (j.f):

3 is a factor of 6 and 15. does this have to do with anything?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \dfrac{1}{2}, \dfrac{1}{12}, \dfrac{1}{30} ,\cdot \cdot \cdot \cdot \cdot \hspace{.33em}\\~\\ \dfrac{1}{2\times 1(2\cdot 1-1)},\dfrac{1}{2\times 2(2\cdot 2-1)}, \dfrac{1}{2\times 3(2\cdot 3-1)} ,\cdot \cdot \cdot \cdot \cdot \hspace{.33em}\\~\\ \normalsize \text{nth term } ~\rightarrow \dfrac{1}{2\times n(2\cdot n-1)} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

for the next term u have to plug in \(n=4\)

hero (hero):

@J.F 1 = (1)(1) 6 = (2)(3) 15 = (3)(5)